# sample.bib

### v1

@@ Abstract
Machine unlearning algorithms aim to efficiently remove data from a model without retraining it from scratch, in order to enforce data privacy, remove corrupted or outdated data, or respect a user’s “right to be forgotten.” Certified machine unlearning is a strong theoretical guarantee that quantifies the extent to which data is erased from the model weights. Most prior works in certified unlearning focus on models trained on convex or strongly convex loss functions, which benefit from convenient convergence guarantees and the existence of global minima. For nonconvex objectives, existing algorithms rely on limiting assumptions and expensive computations that hinder practical implementations. In this work, we propose a simple first-order algorithm for unlearning on general nonconvex loss functions which unlearns by “rewinding” to an earlier step during the learning process and then performs gradient descent on the loss function of the retained data points. Our algorithm is black-box, in that it can be directly applied to models pretrained with vanilla gradient descent with no prior consideration of unlearning. We prove $(\epsilon,\delta)$ certified unlearning and performance guarantees that establish the privacy-utility-complexity tradeoff of our algorithm, with special consideration for nonconvex functions that satisfy the Polyak-Lojasiewicz inequality.
@@ 1 Introduction
Machine unlearning, or data deletion from models, refers to the problem of removing the influence of some data from a trained model without the computational expenses of completely retraining it from scratch [caoOGMUL]. This research direction has become highly relevant in the last few years, due to increasing concern about user privacy and data security as well as the growing cost of retraining massive deep learning models on constantly updated datasets. For example, recent legislation that protects a user’s “Right to be Forgotten,” including the European Union’s General Data Protection Regulation (GDPR), the California Consumer Privacy Act (CCPA), and the Canadian Consumer Privacy Protection Act (CPPA), mandate that users be allowed to request removal of their personal data, which may be stored in databases or memorized by models [sekharirememberwhatyouwant]. In addition, machine unlearning has practical implications for removing the influence of corrupted, outdated, or mislabeled data [nguyen2022surveymachineunlearning] [kurmanji2023towards].
@@
The general goal of machine unlearning algorithms is to yield a model that resembles the model obtained from a full retrain on the updated dataset after data is removed. This requirement is formalized in the concept of “certified unlearning,” a strong theoretical guarantee inspired by differential privacy that probabilistically bounds the difference between the model weights returned by the unlearning and retraining algorithms [Guo]. However, an algorithm only satisfying certified unlearning is not sufficient, as many such algorithms may still be impractical to use. For example, a trivial unlearning algorithm retrains the model from scratch on the retained dataset, but this would provide no efficiency gain. On the other hand, unlearning is also provably satisfied if the learning and unlearning algorithm both output weights randomly sampled from the same Gaussian distribution, but this would yield a poorly performing model. Therefore, an ideal algorithm optimally balances data deletion, model accuracy, and computation, also known as the “privacy-utility-efficiency” tradeoff. Moreover, because training from scratch is computationally expensive in the machine unlearning setting, we desire black-box unlearning algorithms, which can be applied to pretrained models and do not require training to complete with the intention of unlearning later.
@@
Most certified unlearning algorithms are designed for convex or strongly convex functions [Guo] [neel21a] [sekharirememberwhatyouwant]. Relaxing the convexity requirement is challenging since nonconvex functions do not have unique global minima. Recently, there have been several works that provide certified unlearning guarantees for nonconvex functions. For example, [zhang2024towards] proposes a single-step Newton update algorithm with convex regularization followed by Gaussian perturbation, inspired by existing second-order unlearning methods such as [Guo] [sekharirememberwhatyouwant]. Similarly, [qiao2024efficientgeneralizablecertifiedunlearning] proposes a quasi-second-order method that exploits Hessian vector products to avoid directly computing the Hessian. Finally, [chien2024langevin] proposes a first-order method in the form of projected noisy gradient descent with Gaussian noise added at every time step. These algorithms are listed in Table 1.
@@
Prior work has focused on achieving theoretical unlearning guarantees in the nonconvex setting; however, practical unlearning algorithms should also be computationally efficient, convenient to use, unbounded by restrictive assumptions, and yield models that perform well. First, [zhang2024towards] and [qiao2024efficientgeneralizablecertifiedunlearning] are both (quasi) second-order methods, but first-order methods are more computationally efficient. Second, [chien2024langevin] [qiao2024efficientgeneralizablecertifiedunlearning] are not black-box, since during the training process, [chien2024langevin] requires injecting Gaussian noise at every step and [qiao2024efficientgeneralizablecertifiedunlearning] requires recording a statistical vector for each data sample at each time step. These “white-box” algorithms require significant changes to standard learning algorithms, which hinders practical implementation. Finally, both [zhang2024towards] and [chien2024langevin] require a bounded feasible parameter set, which they leverage to provide a loose bound on the distance between the retraining and unlearning outputs dependent on the parameter set radius $R$. However, this yields excessive noise requirements, with the standard deviation $\sigma$ scaling linearly or polynomially with $R$, translating to degraded model performance. See Table 1 for a summary of these results and Section 3 for a more in-depth discussion of $R$ and comparison with prior literature.
Algorithm Loss Function Method Black-box? Bounded $||\theta||\leq R?$
Newton Step [Guo] Strongly convex Second-order No No
Descent-to-delete (D2D) [neel21a] Strongly convex First-order Yes No
Langevin unlearning [chien2024langevin] Nonconvex First-order No Yes
Constrained Newton Step [zhang2024towards] Nonconvex Second-order Yes Yes
Hessian-free Unlearning [qiao2024efficientgeneralizablecertifiedunlearning] Nonconvex Quasi-second-order No No
Our Work (R2D) Nonconvex First-order Yes No
Table 1: Comparison of certified unlearning algorithms for convex and nonconvex functions. The model parameters are denoted by $\theta$.
@@
In this work, we introduce “rewind-to-delete” (R2D), a first-order, black-box, certified unlearning algorithm for general nonconvex functions that does not depend on the parameter set radius $R$. Our learning algorithm consists of vanilla gradient descent steps on the loss function of the dataset followed by Gaussian perturbation. To remove some data from the model, our unlearning algorithm “rewinds” to a model checkpoint from earlier in the original training trajectory, performs additional gradient descent steps on the new loss function, and again adds Gaussian perturbation to the weights. The checkpoint can be saved during the training period, or it can be computed post hoc from a pretrained model via the proximal point method. We show that this algorithm yields $(\epsilon,\delta)$ certified unlearning and provide theoretical guarantees that explicitly address the “privacy-utility-efficiency” tradeoff [liutrilemma] of our unlearning algorithm. Our analysis relies on simple assumptions and does not require a bounded parameter set or projection. In addition, the algorithm is black-box, as it can be applied to a pretrained model as long as the model was trained with vanilla gradient descent.
@@
We also analyze the special case of nonconvex loss functions that satisfy the Polyak-Łojasiewicz (PL) inequality. The PL inequality is a weak condition that guarantees the existence of a connected basin of global minima, to which gradient descent converges at a linear rate [karimiPL]. These functions are especially relevant to deep learning due to the recent discovery that overparameterized neural networks locally satisfy the PL condition in neighborhoods around initialization [LIUPLoverparametrized].
@@
  • We develop a novel ($\epsilon$, $\delta$) certified unlearning algorithm for general nonconvex loss functions. To the best of our knowledge, this is the first nonconvex certified unlearning algorithm that is first-order and black-box.
  • We provide theoretical unlearning guarantees that demonstrate the privacy-utility-efficiency tradeoff of our algorithm, allowing for controllable noise at the expense of privacy or computation. Our analysis requires significantly fewer assumptions than prior work; notably, we do not require that the parameter set is uniformly bounded.
  • We provide utility guarantees of our algorithm, and we show that for the special case of PL loss functions, we obtain linear convergence guarantees and generalization guarantees.
@@ 1.1 Related Work
Differential privacy [dwork2006] is a well-established framework designed to protect individual privacy by ensuring that the inclusion or exclusion of any single data point in a dataset does not significantly affect the output of data analysis or modeling, limiting information leakage about any individual within the dataset. Specifically, a differentially private learning algorithm yields a model trained on some dataset that is probabilistically indistinguishable from a model trained on the same dataset with a data sample removed or replaced. The concept of $(\epsilon,\delta)$-privacy quantifies the strength of this privacy guarantee in terms of the privacy loss, $\epsilon$, and the probability of a privacy breach, $\delta$ [dwork2014algorithmic]. This is typically achieved during or after training by adding controlled noise to the data, model weights [wublackboxdp] [zhangDPijcai2017p548], gradients [abadiDP] [zhangDPijcai2017p548], or objective function [chaudhuri11aDP], in order to mask information about any one sample in the dataset. However, greater noise typically corresponds to worse model performance, leading to a trade-off between utility and privacy. The theory and techniques of differential privacy provide a natural starting point for the rigorous analysis of unlearning algorithms.
@@
As observed in [wublackboxdp], “white-box” differentially private algorithms, which require code changes at every training step to inject noise, are challenging to deploy in the real world due to additional development and runtime overhead incurred from implementing a nonstandard learning algorithm. Rather than adding noise at each iteration, [wublackboxdp] and [zhangDPijcai2017p548] propose only perturbing the output of vanilla stochastic gradient descent, an approach which is easier to integrate into realistic development pipelines. In similar fashion, our proposed black-box unlearning algorithm can be implemented without any special steps during learning with gradient descent, and we also mask the difference in algorithm outputs with noise at the end of training. The difference between our approach and [wublackboxdp] [zhangDPijcai2017p548], however, is that our approach can accommodate the nonconvex case, leveraging model checkpointing to control the distance from the retraining trajectory.
Certified unlearning.
@@
The term “machine unlearning” was first coined by [caoOGMUL] to describe a deterministic data deletion algorithm, which has limited application to general optimization problems. In the following years, techniques have been developed for “exact unlearning,” which exactly removes the influence of data, and “approximate unlearning,” which yields a model that is approximately close to the retrained model with some determined precision [xusurvey]. Our work focuses on the latter. Both [ginartmakingAI] and [Guo] propose a probabilistic notion of approximate unlearning, where the unlearning and retraining outputs must be close in distribution. Inspired by differential privacy, [Guo] introduces the definition of certified $(\epsilon,\delta)$ unlearning that we use in this work. Like differentially private algorithms, certified unlearning algorithms typically require injected noise to the weights or objective function, which can degrade model performance. However, unlearning algorithms are also designed to reduce computation, leading to a three-way trade-off between privacy, utility, and complexity.
@@
Certified unlearning has been studied for a variety of specific settings, including linear and logistic models [Guo] [izzo21a], graph neural networks [chien2022certifiedgraphunlearning], and minimax models [liu2023certified], as well as convex models [sekharirememberwhatyouwant] [neel21a] [chien2024stochastic] [suriyakumar2022algorithms] and nonconvex models [chien2024langevin] [qiao2024efficientgeneralizablecertifiedunlearning] [zhang2024towards]. Specifically, our work is closely inspired by [neel21a], which proposes “descent-to-delete,” a first-order unlearning algorithm for strongly convex functions that unlearns by fine-tuning with gradient descent iterates on the loss function of the retained samples. The algorithms for general models can be categorized as first-order methods that only require access to the function gradients [neel21a] [chien2024stochastic] or second-order methods that leverage information from the Hessian to approximate the model weights that would result from retraining [sekharirememberwhatyouwant] [suriyakumar2022algorithms] [zhang2024towards] [qiao2024efficientgeneralizablecertifiedunlearning].
Nonconvex unlearning.
@@
In addition to the works mentioned above, there are many approximate unlearning algorithms designed for nonconvex functions such as neural networks that have weaker theoretical guarantees. For example, [bui2024newtonsmethodunlearnneural] proposes a “weak unlearning” algorithm, which considers indistinguishability with respect to model output space instead of model parameter space. This is a weaker guarantee because the model weights may still retain information about the unlearned data and be susceptible to membership inference attacks. Another popular algorithm for neural networks is SCRUB [kurmanji2023towards], a gradient-based algorithm that balances maximizing error on the forget set and maintaining performance on the retain set. An extension of SCRUB, SCRUB+Rewind, “rewinds” the algorithm to a point where the error on the forget set is “just high enough,” so as to impede membership inference attacks. Finally, although [golatkar1] [golatkar2] propose unlearning algorithms for deep neural networks, they only provide a general upper bound on the amount of information retained in the weights rather than a strict certified unlearning guarantee.
@@
Additional approaches include subtracting out the impact of the forget datapoints in each batch of gradient descent [Graves_Nagisetty_Ganesh_2021], gradient ascent on the loss function of the forget set [jang-etal-2023-knowledge], and retraining the last layers of the neural network on the retain set [goel]. Ultimately, while the ideas of checkpointing, gradient ascent, and “rewinding” have been considered in other machine unlearning works, our algorithm combines these elements in a novel manner to obtain strong theoretical guarantees that prior algorithms lack.

### v2

# Rewind-to-Delete: Certified Machine Unlearning for Nonconvex Functions

@@ Abstract
Machine unlearning algorithms aim to efficiently remove data from a model without retraining it from scratch, in order to remove corrupted or outdated data or respect a user’s “right to be forgotten.” Certified machine unlearning is a strong theoretical guarantee based on differential privacy that quantifies the extent to which an algorithm erases data from the model weights. In contrast to existing works in certified unlearning for convex or strongly convex loss functions, or nonconvex objectives with limiting assumptions, we propose the first, first-order, black-box (i.e., can be applied to models pretrained with vanilla gradient descent) algorithm for unlearning on general nonconvex loss functions, which unlearns by “rewinding” to an earlier step during the learning process before performing gradient descent on the loss function of the retained data points. We prove $(\epsilon,\delta)$ certified unlearning and performance guarantees that establish the privacy-utility-complexity tradeoff of our algorithm, and we prove generalization guarantees for nonconvex functions that satisfy the Polyak-Lojasiewicz inequality. Finally, we implement our algorithm under a new experimental framework that more accurately reflects real-world use cases for preserving user privacy.
Machine Learning, ICML
@@ 1 Introduction
Machine unlearning, or data deletion from models, refers to the problem of removing the influence of some data from a trained model without the computational expenses of completely retraining it from scratch (Cao & Yang, 2015). This research direction has become highly relevant in the last few years, due to increasing concern about user privacy and data security as well as the growing cost of retraining massive deep learning models on constantly updated datasets. For example, recent legislation that protects a user’s “Right to be Forgotten,” including the European Union’s General Data Protection Regulation (GDPR), the California Consumer Privacy Act (CCPA), and the Canadian Consumer Privacy Protection Act (CPPA), mandate that users be allowed to request removal of their personal data, which may be stored in databases or memorized by models (Sekhari et al., 2021). In addition, machine unlearning has practical implications for removing the influence of corrupted, outdated, or mislabeled data (Kurmanji et al., 2023; Nguyen et al., 2024).
@@
The typical goal of machine unlearning algorithms is to yield a model that resembles the model obtained from a full retrain on the updated dataset after data is removed. This requirement is formalized in the concept of “certified unlearning,” a strong theoretical guarantee motivated by differential privacy that probabilistically bounds the difference between the model weights returned by the unlearning and retraining algorithms (Guo et al., 2019). However, algorithms satisfying certified unlearning need to also be practical. For example, a trivial unlearning algorithm retrains the model from scratch on the retained dataset, but this would provide no efficiency gain. On the other hand, unlearning is also provably satisfied if the learning and unlearning algorithm both output weights randomly sampled from the same Gaussian distribution, but this would yield a poorly performing model. Therefore, an ideal algorithm optimally balances data deletion, model accuracy, and computation, also known as the “privacy-utility-efficiency” tradeoff (Liu et al., 2024). Moreover, because training from scratch is computationally expensive in the machine unlearning setting, we desire black-box unlearning algorithms, which can be applied to pretrained models and do not require training with the intention of unlearning later.
@@
Most certified unlearning algorithms are designed for convex or strongly convex functions (Guo et al., 2019; Neel et al., 2021; Sekhari et al., 2021). Relaxing the convexity requirement is challenging since nonconvex functions do not have unique global minima. Recently, there have been several works that provide certified unlearning guarantees for nonconvex functions. For example, (Zhang et al., 2024) proposes a single-step Newton update algorithm with convex regularization followed by Gaussian perturbation, inspired by existing second-order unlearning methods such as (Guo et al., 2019; Sekhari et al., 2021). Similarly, (Qiao et al., 2024) proposes a quasi-second-order method that exploits Hessian vector products to avoid directly computing the Hessian. Finally, (Chien et al., 2024a) proposes a first-order method in the form of projected noisy gradient descent with Gaussian noise added at every step.
@@
Prior work has focused on achieving theoretical unlearning guarantees in the nonconvex setting; however, practical unlearning algorithms should also be computationally efficient and convenient to use. First, (Zhang et al., 2024) and (Qiao et al., 2024) are both (quasi) second-order methods, but first-order methods that only require computing the gradient are more computationally efficient and require less storage. Second, (Chien et al., 2024a; Qiao et al., 2024) are not black-box, since during the training process, (Chien et al., 2024a) requires injecting Gaussian noise at every step and (Qiao et al., 2024) requires recording a statistical vector for each data sample at each time step. These “white-box” algorithms require significant changes to standard learning algorithms, which hinders easy implementation. Table 1 summarizes these results and Appendix C provides a more thorough comparison with prior work.
@@
In this work, we introduce “rewind-to-delete” (R2D), a first-order, black-box, certified unlearning algorithm for general nonconvex functions. Our learning algorithm consists of vanilla gradient descent steps on the loss function of the dataset followed by Gaussian perturbation. To remove some data from the model, our unlearning algorithm “rewinds” to a model checkpoint from earlier in the original training trajectory, performs additional gradient descent steps on the new loss function, and again adds Gaussian perturbation to the weights. The checkpoint can be saved during the training period (which is standard practice), or it can be computed post hoc from a pretrained model via the proximal point method. We prove $(\epsilon,\delta)$ certified unlearning for our algorithm and provide theoretical guarantees that explicitly address the “privacy-utility-efficiency” tradeoff of our unlearning algorithm. Our algorithm is easy to implement and relies on simple assumptions. In addition, the algorithm is black-box, as it can be applied to a pretrained model as long as the model was trained with vanilla gradient descent.
Table 1: Comparison of certified unlearning algorithms for convex and nonconvex functions. $d$ is the dimension of the model parameters and $n$ is the size of the training dataset.
Algorithm Loss        Function Method Storage Black-box?
Newton Step111(Guo et al., 2019) Strongly convex Second-order $O(d^{2})$$\times$
Descent-to-delete 222(Neel et al., 2021) Strongly convex First-order $O(d)$ $\surd$
Langevin unlearning333(Chien et al., 2024a) Nonconvex First-order $O(d)$$\times$
Constrained Newton Step444(Zhang et al., 2024) Nonconvex Second-order $O(d^{2})$ $\surd$
Hessian-Free Unlearning 555(Qiao et al., 2024) Nonconvex Quasi-second-order $O(nd)$$\times$
Our Work (R2D) Nonconvex First-order $O(d)$ $\surd$
@@
We also analyze the case of nonconvex functions that satisfy the Polyak-Łojasiewicz (PL) inequality. The PL inequality is a weak condition that guarantees the existence of a connected basin of global minima, to which gradient descent converges at a linear rate (Karimi et al., 2016). This property allows us to derive empirical risk convergence and generalization guarantees, despite nonconvexity. PL functions are highly relevant to deep learning because overparameterized neural networks locally satisfy the PL condition in neighborhoods around initialization (Liu et al., 2022).
Finally, we empirically demonstrate the privacy-utility-efficiency tradeoff of our algorithm and its superior performance compared to existing certified unlearning algorithms for nonconvex functions. We use real-world medical or facial data, with each dataset derived from a collection of users. In contrast to most unlearning experiments, which either select unlearned samples uniformly at random or from a specific class, we train a neural network model to learn some global characteristics about a dataset, and unlearn data from a subset of the users, thereby simulating realistic unlearning requests and their impact on model utility. This results in an unlearned dataset that is not i.i.d. to the training or test set. To assess unlearning empirically, we conduct a membership inference attack (MIA) (Shokri et al., 2016) comparing the unlearned model output on the unlearned dataset and an out-of-distribution dataset, constructed separately from data and users not present in the training data. Comparing with baseline methods, on the medical dataset we outperform (Zhang et al., 2024) by 9-24 accuracy percentage points using 41%-80% rewinding, and on the facial dataset we outperform (Zhang et al., 2024) by 23-36 accuracy percentage points and 3%-5% on unlearning, using 51%-75% rewinding, and we are comparable in compute time. [^6] See Table 8 and 9 in the Appendix for numerical details. Compared to the white-box algorithm (Qiao et al., 2024), we are competitive in accuracy and unlearning and 70-104 times faster during training, despite using more steps. Ultimately, our first-order method enjoys the benefits of minimal computational requirements, easy off-the-shelf implementation, and competitive performance.
@@
  • We develop the first ($\epsilon$, $\delta$) certified unlearning algorithm for nonconvex loss functions that is first-order and black-box.
  • We prove theoretical unlearning guarantees that demonstrate the privacy-utility-efficiency tradeoff of our algorithm, allowing for controllable noise at the expense of privacy or computation. For the special case of PL loss functions, we obtain linear convergence and generalization guarantees.
  • We conduct experiments demonstrating the efficacy and utility of our algorithm under a novel experimental framework that better reflects real-world unlearning scenarios for protecting user privacy.
@@ 1.1 Related Work
Differential privacy. Differential privacy (DP) (Dwork et al., 2006) is a well-established framework designed to protect individual privacy by ensuring that the inclusion or exclusion of any single data point in a dataset does not significantly affect the output of data analysis or modeling, limiting information leakage about any individual within the dataset. Specifically, a differentially private learning algorithm yields a model trained on some dataset that is probabilistically indistinguishable from a model trained on the same dataset with a data sample removed or replaced. The concept of $(\epsilon,\delta)$-privacy quantifies the strength of this privacy guarantee in terms of the privacy loss, $\epsilon$, and the probability of a privacy breach, $\delta$ (Dwork & Roth, 2014). This privacy can be injected during or after training by adding controlled noise to the data, model weights (Wu et al., 2017; Zhang et al., 2017), gradients (Abadi et al., 2016; Zhang et al., 2017), or objective function (Chaudhuri et al., 2011), in order to mask information about any one sample in the dataset. However, greater noise typically corresponds to worse model performance, leading to a trade-off between utility and privacy. The theory and techniques of differential privacy provide a natural starting point for the rigorous analysis of unlearning algorithms.
@@
As observed in (Wu et al., 2017), “white-box” differentially private algorithms, which require code changes to inject noise at every training step, are challenging to deploy in the real world due to additional development and runtime overhead incurred from implementing a nonstandard learning algorithm. Rather than adding noise at each iteration, (Wu et al., 2017) and (Zhang et al., 2017) propose DP algorithms that only perturb the output after training, an approach which is easier to integrate into real-world development pipelines. In similar fashion, our proposed black-box unlearning algorithm can be implemented without any special steps during learning with gradient descent. We also do not inject noise at each iteration, only perturbing at the end of training. The difference between our approach and (Wu et al., 2017; Zhang et al., 2017), however, is that our approach can accommodate the nonconvex case, leveraging model checkpointing to control the distance from the retraining trajectory.
@@
Certified unlearning. The term “machine unlearning” was first coined by (Cao & Yang, 2015) to describe a deterministic data deletion algorithm, which has limited application to general optimization problems. In the following years, techniques have been developed for “exact unlearning,” which exactly removes the influence of data, and “approximate unlearning,” which yields a model that is approximately close to the retrained model with some determined precision (Xu et al., 2023). Our work focuses on the latter. Both (Ginart et al., 2019) and (Guo et al., 2019) introduce a probabilistic notion of approximate unlearning, where the unlearning and retraining outputs must be close in distribution. Inspired by differential privacy, (Guo et al., 2019) introduces the definition of certified $(\epsilon,\delta)$ unlearning used in this work. Like DP algorithms, certified unlearning algorithms typically require injected noise to the weights or objective function, which can degrade model performance. Moreover, unlearning algorithms are also designed to reduce computation, leading to a three-way trade-off between privacy, utility, and complexity.
@@
Certified unlearning has been studied for a variety of settings, including linear and logistic models (Guo et al., 2019; Izzo et al., 2021), graph neural networks (Chien et al., 2022), minimax models (Liu et al., 2023), and the federated learning setting (Fraboni et al., 2024), as well as convex models (Sekhari et al., 2021; Neel et al., 2021; Suriyakumar & Wilson, 2022; Chien et al., 2024b) and nonconvex models (Chien et al., 2024a; Qiao et al., 2024; Zhang et al., 2024). These algorithms can be categorized as first-order methods that only require access to the function gradients (Neel et al., 2021; Chien et al., 2024b) or second-order methods that leverage information from the Hessian to approximate the model weights that would result from retraining (Sekhari et al., 2021; Suriyakumar & Wilson, 2022; Zhang et al., 2024; Qiao et al., 2024). Our work is inspired by the “descent-to-delete” algorithm (Neel et al., 2021), a first-order unlearning algorithm for strongly convex functions that unlearns by fine-tuning with gradient descent iterates on the loss function of the retained samples.
@@
Nonconvex unlearning. There are also many approximate unlearning algorithms for nonconvex functions that rely on heuristics or weaker theoretical guarantees. For example, (Bui et al., 2024) proposes a “weak unlearning” algorithm, which considers indistinguishability with respect to model output space instead of model parameter space. This is a weaker guarantee because the model weights may still retain information about the unlearned data and be susceptible to membership inference attacks. Another popular algorithm for neural networks is SCRUB (Kurmanji et al., 2023), a gradient-based algorithm that balances maximizing error on the unlearned data and maintaining performance on the retained data. An extension of SCRUB, SCRUB+Rewind, “rewinds” the algorithm to a point where the error on the unlearned data is “just high enough,” so as to impede membership inference attacks. Finally, although (Golatkar et al., 2019, 2020) propose unlearning algorithms for deep neural networks, they only provide a general upper bound on the amount of information retained in the weights rather than a strict certified unlearning guarantee.
@@
Additional approaches include subtracting out the impact of the unlearned data in each batch of gradient descent (Graves et al., 2021), gradient ascent on the loss function of the unlearned data (Jang et al., 2023), and retraining the last layers of the neural network on the retained data (Goel et al., 2022). Ultimately, while the ideas of checkpointing, gradient ascent, and “rewinding” have been considered in other machine unlearning works, our algorithm combines these elements in a novel manner to obtain strong theoretical guarantees that prior algorithms lack.